# Diophantine approximation with sums of two squares

### v3

@@ Abstract.
For any given positive definite binary quadratic form $Q$ with integer coefficients, we establish two results on Diophantine approximation with integers represented by $Q$. Firstly, we show that for every irrational number $\alpha$, there exist infinitely many positive integers $n$ represented by $Q$ and satisfying $\alpha n||<n^{-(1/2-\varepsilon)}$ for any fixed but arbitrarily small $\varepsilon>0$. This is an easy consequence of a result by Cook on small fractional parts of diagonal quadratic forms. Secondly, we give a quantitative version with a lower bound of this result when the exponent $1/2-\varepsilon$ is replaced by any fixed $\gamma<3/7$. To this end, we use the Voronoi summation formula and a bound for bilinear forms with Kloosterman sums to fixed moduli by Kerr, Shparlinski, Wu and Xi.
@@ 1. Introduction and main results
Throughout this article, we denote by $z||$ the distance of $z\in\mathbb{R}$ to the nearest integer. As usual, $\varepsilon$ stands for an arbitrarily small positive number. Moreover, we write
@@
A basic theorem in Diophantine approximation is the Dirichlet approximation theorem, which implies that for every irrational number $\alpha$ there are infinitely many positive integers $n$ such that $\alpha n||<n^{-1}$, where $x||$ is the distance of $x$ from the nearest integer. Beyond this classical result, mathematicians have explored obtaining good approximations when $n$ is restricted to sparse subsets of the set of positive integers. Numerous results exist in this direction. Here we mention only a few of them. For example, when $n$ varies over the set of primes, the best known approximation result says that $\alpha n||<n^{-1/3+\varepsilon}$ holds for infinitely many primes $n$. This is due to Matomäki [11]. In the case when $n$ varies over the set of square-free positive integers, it was established by Heath-Brown [7] that $\alpha n||<n^{-2/3+\varepsilon}$ holds for infinitely many squarefree $n\in\mathbb{N}$. In this article, we consider the case where $n$ varies over positive integers which are represented by any given positive definite binary quadratic form with integer coefficients.
Let $Q(x,y)=a_{1}x^{2}+b_{1}xy+c_{1}y^{2}$ be a positive definite integral binary quadratic form (PBQF). In particular, the discriminant $b_{1}^{2}-4a_{1}c_{1}$ of $Q$ is negative. We set
@@
Theorem 2 (Main Theorem, [6]).
Let
@@
Theorem 3 (Theorem 1, [5]).
Let $k\geq 2$ and $K=2^{k-1}.$ Let $f(\vec{x})=\theta_{1}x_{1}^{k}+\theta_{2}x_{2}^{k}+\cdots+\theta_{s}x_{s}^{k}$ be an additive form of degree $k$, in $s$ variables with $1\leq s\leq K$, where $\theta_{i}$ are any real numbers and $\vec{x}=(x_{1},x_{2},\cdots,x_{s}).$ Then for every $\epsilon>0,N>1$, there are integers $x_{1},x_{2},\ldots,x_{s}$, not all zero, such that
[…]
@@
[…]
for another constant $D_{2}(\varepsilon,Q)>0$ since $Q$ is positive definite. Note that $\alpha n||<D_{1}(\varepsilon,Q)N^{-1+\varepsilon}$ does not hold for arbitrarily large $N$ because $\alpha$ is irrational. Hence, there is an infinite sequence of natural numbers $n$ such that
$\alpha n||<D_{2}(\varepsilon,Q)n^{-1/2+\varepsilon/2},$
@@
Danicic [6] used lattices to establish Theorem 2, and Cook [5] used bounds for quadratic Weyl sums to establish Theorem 3. Both methods give only the existence of infinitely many good approximations but no bounds on the number of them. The main part of this article consists of establishing a lower bound which holds for certain ranges of $n$ depending on Diophantine approximations of $\alpha$. This comes at the cost of a weaker exponent of $3/7$ in place of $1/2$, which still beats the exponent $1/3$ coming from Theorem 2. Our method uses the Voronoi summation formula and bounds for bilinear sums with Kloosterman sums to a fixed modulus due to Kerr, Shparlinski, Wu and Xi [10]. We will deduce our result from the following asymptotic formula for the number of representations by $Q$ in unions of residue classes, which is of independent interest.
@@
Let $q\in\mathbb{N}$ and $a\in\mathbb{Z}$ such that $(2\Delta a,q)=1$. Let $\varepsilon>0$ and $L=q^{\beta},X=Lq=q^{1+\beta}$ with $2/5+3\varepsilon\leq\beta<1$. Let $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ be a smooth function which is compactly supported in $[X,2X]$ with $w(x)=1$ in $[5X/4,7X/4]$ such that for any integer $j\geq 1$, the $j^{th}$ derivative satisfies
$w^{(j)}(x)\ll_{j}X^{-j}.$ (1.3)
@@
$S:=\sum\limits_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum\limits_{\begin{subarray}{c}n\\ na\equiv b\bmod{q}\end{subarray}}r_{Q}(n)w(n),$ (1.4)
where $r_{Q}(n)$ is the number of integral representations of $n$ by the PBQF $Q(x,y)$. Then for sufficiently large $X$, we have
$S=\frac{4\pi\varphi(q)L}{\Delta^{1/2}q^{2}}\prod_{\begin{subarray}{c}p|q\\ p\text{ prime }\end{subarray}}\left(1-\frac{G_{Q}(p,1)}{p^{2}}\right)\int\limits_{0}^{\infty}w(x)dx+O_{\varepsilon}\left(\frac{XL}{q}q^{-\varepsilon}\right),$ (1.5)
where
@@
$\frac{XL}{q\log^{2}q}\ll S\ll\frac{XL}{q}.$
We will use Theorem 4 above to deduce the following lower bound for Diophantine approximations with integers represented by $Q$, which is close to the expected bound.
@@
[…]
@@
We note that (12/5)/(1+2/5)=3/7(1-2/5)/(1+2/5)=3/7, and thus Corollary 1 recovers Theorem 1 with a weaker exponent of $3/7$ in place of $1/2$, but coming with a lower bound for the number of $n$’s satisfying $n\alpha||<n^{-3/7+\varepsilon}$.
@@
They started in a similar way as we do here, converting the Diophantine condition $n\alpha||<\delta$ into a condition of $n$ lying in a union of residue classes modulo $q$. Then they proceeded by detecting the squarefreeness of $n$ using the convolution identity $\mu^{2}(n)=\sum\limits_{t^{2}|n}\mu(t)$ and the congruence condition using the orthogonality relations for Dirichlet characters. After isolating a main term, they were left with character sums which were treated using Hölder’s inequality and moment bounds for character sums. This yielded an exponent of $1/2-\varepsilon$ for the relevant Diophantine approximation problem. To handle Diophantine approximation with denominators represented by sums of two squares following their approach, one may use the convolution identiy $r_{2}(n)=4\sum\limits_{d|n}\chi_{4}(d)$ with $r_{2}(n)$ being the number of representations of $n$ as a sum of two squares and $\chi_{4}$ the non-trivial character modulo 4 . However, here we have a divisor condition of $d|n$ in place of $t^{2}|n$. As a consequence (even if one uses the hyperbola trick), one gets only an exponent of $1/3-\varepsilon$ instead of $1/2-\varepsilon$ in the Diophantine approximation problem for sums of two squares. We also note that their method does not extend to general quadratic forms $Q$ because of the lack of a simple formula for the number of representations of $n$ in the form $n=Q(x,y)$.
@@ 2. Preliminaries
where $A$ is an integral symmetric $2\times 2$ matrix with even diagonal elements, and $(x,y)^{T}$ is the column vector of the indeterminates. Then the adjoint matrix $A^{\dagger}$ of $A$ is defined by $A^{\dagger}A=\Delta I_{2}$, where $\Delta$ is the determinant of the matrix $A$ and $I_{2}$ is the $2\times 2$ identity matrix. Clearly, $A^{\dagger}$ has also integer entries.
Our main tools are the Voronoi summation formula for quadratic forms and bounds for bilinear sums with Kloosterman sums to a fixed modulus. To state the Voronoi summation formula, let us first set up some notations. Let $\mathbb{P}$ be the set of primes. For a prime $p$ and integer $n\not=0$, let $\operatorname{ord}_{p}(n)$ be the largest $k\in\mathbb{N}_{0}$ such that $p^{k}$ divides $n$. For given positive integer $q$, we define
$\delta_{0}=\delta_{0}(q):=\prod_{\begin{subarray}{c}p\in\mathbb{P}\\ \operatorname{ord}_{p}(\Delta)\leq\operatorname{ord}_{p}(q)\end{subarray}}p^{\operatorname{ord}_{p}(\Delta)}$ (2.1)
and
$\delta_{1}=\delta_{1}(q):=\prod_{\begin{subarray}{c}p\in\mathbb{P}\\ \operatorname{ord}_{p}(\Delta)>\operatorname{ord}_{p}(q)\end{subarray}}p^{\operatorname{ord}_{p}(q)}.$ (2.2)
Clearly, one can write $\delta:=(\Delta,q)=\delta_{0}\delta_{1}$. Here, as in the following, $(\Delta,q)$ denotes the gcd of $\Delta$ and $q$. Let $\Delta_{0}=\Delta/\delta_{0},\ \ q_{1}=q/\delta_{1}$. We note that $(\delta_{1},q_{1})=(\Delta_{0},q)=(\delta_{0},\delta_{1})=1$. Also, for given coprime integers $q$ and $h$, we define the Gauss sum associated to the form $Q$ as
@@
More generally, for any vectors $\vec{x},\vec{a}\in\mathbb{Z}^{2}$, we define
$G_{Q}(q,h,\vec{a}):=\sum_{\vec{x}\bmod q}e_{q}(hQ(\vec{x}+\vec{a}\cdot\vec{x}),$ (2.4)
where $\vec{a}\cdot\vec{x}$ denotes the standard inner product of vectors. The following lemma gives an evaluation of $G_{Q}(q,h,\vec{a})$.
@@
Let $\Delta,q,h,\delta_{0},\delta_{1},\Delta_{0},q_{1}$ be defined as above and let $\bar{\Delta}_{0}$ be the multiplicative inverse of $\Delta$ modulo $q_{1}$ such that $\delta_{1}|\bar{\Delta}_{0}$. Let $g$ be an integer such that
$gq_{1}\equiv 1(\bmod\delta_{1}).$ (2.5)
(i) Then for any vector $\vec{a}\in\mathbb{Z}^{2}$, we have
$G_{Q}(q,h,\vec{a})=E_{\delta_{0}}(A^{\dagger}\vec{a})e_{q_{1}}(-\bar{h}\bar{\Delta}_{0}\delta^{-1}Q^{\dagger}(\vec{a}))\sum_{\vec{x}(\bmod q)}e_{q}(hQ(\vec{x}))e_{\delta_{1}}(\vec{x}\cdot g\vec{a}),$ (2.6)
where $Q^{\dagger}$ is the quadratic form associated to the adjoint matrix $A^{\dagger}$ and
$E_{m}(\vec{x})=\begin{cases}1\ \ \text{ if }\vec{x}\equiv\vec{0}(\bmod{m}),\\ 0\ \ \text{ otherwise.}\end{cases}$ (2.7)
(ii) For $q$ odd, we have
@@
$N(q,Q):=\sharp\{\vec{x}(\bmod q):A\vec{x}\equiv\vec{0}(\bmod q)\}.$ (2.9)
@@
See [3, Theorem 1] or, [9, Theorem 1]. ∎
@@
For given positive integer $q$ and any integer $h$ with $(q,h)=1$, let $G_{Q}(q,h)$ be defined by (2.3). Then, we have
@@
$b_{i}=\frac{\delta_{0}}{(\delta_{0},s_{i})},\ B=(b_{i}\delta_{ij}),\ A^{*}=\delta_{0}^{-1}BV^{T}A^{\dagger}VB.$ (2.11)
@@
$\tilde{r}_{Q^{*}}(n,q,h):=q^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\ \sum_{\vec{y}(\bmod q)}e_{q}(hQ(\vec{y}))e_{\delta_{1}}(\vec{y}\cdot gVB\vec{x}),$ (2.12)
where $\delta_{0}=\delta_{0}(q),\delta_{1}=\delta_{1}(q),g$ be given in (2.1), (2.2), (2.5), respectively. Now we are ready to state the Voronoi summation formula for the quadratic form $Q$.
@@
Let $f(n)$ be a continuously differentiable function on $\mathbb{R}$ with compact support in $\mathbb{R}_{\geq 0}$. Then for $(h,q)=1$ with $q\geq 1$, we have
$\sum_{n}r_{Q}(n)e_{q}(nh)f(n)=2\pi\Delta^{-1/2}q^{-2}G_{Q}(q,h)\int\limits_{0}^{\infty}f(x)dx$ (2.13)
$+2\pi(q\sqrt{\Delta})^{-1}\sum_{n}\tilde{r}_{Q^{*}}(n,q,h)e_{q_{1}}(-\bar{h}\bar{\Delta}_{0}\delta^{-1}n)\int\limits_{0}^{\infty}J_{0}\left(4\pi\frac{\sqrt{nx/\Delta_{0}}}{q}\right)f(x)dx,$ (2.14)
where $\tilde{r}_{Q^{*}}(n,q,h)$ is given in (2.12), $\bar{\Delta}_{0}$ is a multiplicative inverse of $\Delta_{0}$ modulo $q_{1}$ and $J_{0}(x)$ is the standard Bessel function.
@@
See the proof of [9, Equation (28)]. ∎
@@
$S(m,n;q):=\sum_{\begin{subarray}{c}1\leq x\leq q-1\\ (x,q)=1\end{subarray}}e\left(\frac{mx+n\bar{x}}{q}\right).$
@@
The following Dirichlet approximation type lemma will allow us to choose the modulus $q$ coprime to $2\Delta$ (in fact coprime to any given non-zero integer). This will simplify our calculations when using Lemma 3.
@@
Fix $d\in\mathbb{N}$ and $\alpha\in\mathbb{R}\setminus\mathbb{Q}$. We aim to find good approximations of $\alpha$ by fractions $b/r$ with $q$ odd and $(r,bd)=1$. Take one of the infinitely many approximations of $\alpha$ by $a/q$ with $(a,q)=1$ and

### v4

@@ Abstract.
For any given positive definite binary quadratic form $Q$ with integer coefficients, we establish two results on Diophantine approximation with integers represented by $Q$. Firstly, we show that for every irrational number $\alpha$, there exist infinitely many positive integers $n$ represented by $Q$ and satisfying $\|\alpha n\|<n^{-(1/2-\varepsilon)}$ for any fixed but arbitrarily small $\varepsilon>0$. This is an easy consequence of a result by Cook on small fractional parts of diagonal quadratic forms. Secondly, we give a quantitative version with a lower bound of this result when the exponent $1/2-\varepsilon$ is replaced by any fixed $\gamma<3/7$. To this end, we use the Voronoi summation formula and a bound for bilinear forms with Kloosterman sums to fixed moduli by Kerr, Shparlinski, Wu and Xi.
@@ 1. Introduction and main results
Throughout this article, we denote by $\|z\|$ the distance of $z\in\mathbb{R}$ to the nearest integer. As usual, $\varepsilon$ stands for an arbitrarily small positive number. Moreover, we write
@@
A basic theorem in Diophantine approximation is the Dirichlet approximation theorem, which implies that for every irrational number $\alpha$ there are infinitely many positive integers $n$ such that $\|\alpha n\|<n^{-1}$, where $\|x\|$ is the distance of $x$ from the nearest integer. Beyond this classical result, mathematicians have explored obtaining good approximations when $n$ is restricted to sparse subsets of the set of positive integers. Numerous results exist in this direction. Here we mention only a few of them. For example, when $n$ varies over the set of primes, the best known approximation result states that $\|\alpha n\|<n^{-1/3+\varepsilon}$ holds for infinitely many primes $n$. This is due to Matomäki [11]. In the case when $n$ varies over the set of square-free positive integers, it was established by Heath-Brown [8] that $\|\alpha n\|<n^{-2/3+\varepsilon}$ holds for infinitely many squarefree $n\in\mathbb{N}$. In this article, we consider the case where $n$ varies over positive integers which are represented by any given positive definite binary quadratic form with integer coefficients.
Let
$Q(x,y)=a_{1}x^{2}+b_{1}xy+c_{1}y^{2}$
be a positive definite integral binary quadratic form (PBQF). In particular, the discriminant $b_{1}^{2}-4a_{1}c_{1}$ of $Q$ is negative. We set
@@
Theorem 2.
[6, Main Theorem] Let
@@
Theorem 3.
[5, Theorem 1] Let $k\geq 2$ and $K=2^{k-1}.$ Let $f(\vec{x})=\theta_{1}x_{1}^{k}+\theta_{2}x_{2}^{k}+\cdots+\theta_{s}x_{s}^{k}$ be an additive form of degree $k$, in $s$ variables with $1\leq s\leq K$, where $\theta_{i}$ are any real numbers and $\vec{x}=(x_{1},x_{2},\cdots,x_{s}).$ Then for every $\epsilon>0,N>1$, there are integers $x_{1},x_{2},\ldots,x_{s}$, not all zero, such that
$\max\limits_{1\leq i\leq s}|x_{i}|\leq N\ \text{ \ and \ }\ \|f(\vec{x})\|<C(\varepsilon,k)N^{-s/K+\varepsilon}.$
@@
$\|\alpha n\|<D_{1}(\varepsilon,Q)N^{-1+\varepsilon}\leq D_{2}(\varepsilon,Q)n^{-1/2+\varepsilon/2}$
for another constant $D_{2}(\varepsilon,Q)>0$ since $Q$ is positive definite. Note that $\|\alpha n\|<D_{1}(\varepsilon,Q)N^{-1+\varepsilon}$ does not hold for arbitrarily large $N$ because $\alpha$ is irrational. Hence, there is an infinite sequence of natural numbers $n$ such that
$\|\alpha n\|<D_{2}(\varepsilon,Q)n^{-1/2+\varepsilon/2},$
@@
Danicic [6] used lattices to establish Theorem 2, and Cook [5] used bounds for quadratic Weyl sums to establish Theorem 3. Both methods give only the existence of infinitely many good approximations but no bounds on the number of them. The main part of this article consists of establishing a lower bound which holds for certain ranges of $n$ depending on good rational approximations of $\alpha$. This comes at the cost of a weaker exponent of $3/7$ in place of $1/2$, which still beats the exponent $1/3$ coming from Theorem 2. Our method uses the Voronoi summation formula and bounds for bilinear sums with Kloosterman sums to a fixed modulus due to Kerr, Shparlinski, Wu and Xi [10]. We will deduce our result from the following asymptotic formula for the number of representations by $Q$ in unions of residue classes, which is of independent interest.
@@
Suppose that $q\in\mathbb{N}$ and $a\in\mathbb{Z}$ such that $(2\Delta a,q)=1$. Let $\varepsilon>0$ and set $L:=q^{\beta}$ and $X:=Lq=q^{1+\beta}$ with $2/5+3\varepsilon\leq\beta<1$. Let $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ be a smooth function which is compactly supported in $[X,2X]$ with $w(x)=1$ in $[5X/4,7X/4]$ such that for any integer $j\geq 1$, the $j^{th}$ derivative satisfies
$w^{(j)}(x)\ll_{j}X^{-j}.$
@@
$S:=\sum\limits_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum\limits_{\begin{subarray}{c}n\in\mathbb{Z}\\ na\equiv b\bmod{q}\end{subarray}}r_{Q}(n)w(n),$ (1.3)
where $r_{Q}(n)$ is the number of integral representations of $n$ by the PBQF $Q(x,y)$. Then we have
$S=\frac{4\pi\varphi(q)L}{\Delta^{1/2}q^{2}}\prod_{\begin{subarray}{c}p|q\\ p\text{ prime }\end{subarray}}\left(1-\frac{G_{Q}(p,1)}{p^{2}}\right)\int\limits_{0}^{\infty}w(x)dx\cdot\left(1+o(1)\right)$
as $q\rightarrow\infty$, where
@@
$\frac{q^{2\beta}}{(\log\log q)^{2}}\ll S\ll q^{2\beta}.$
We will use Theorem 4 above to deduce the following lower bound for the number of $n$’s represented by $Q$ satisfying $\|n\alpha\|<C_{1}n^{-\gamma}$ for some constant $C_{1}$ and exponent $\gamma$ in a suitable range. This lower bound is close to the expected bound.
@@
$\sharp\{n\in\mathcal{A}_{Q}:n\leq 2X,\ \|n\alpha\|<C_{1}n^{-\gamma}\}\gg_{\varepsilon,\eta}X^{1-\gamma-C_{2}/(\log\log X)}$
@@
We note that (12/5)/(1+2/5)=3/7(1-2/5)/(1+2/5)=3/7, and thus Corollary 1 recovers Theorem 1 with a weaker exponent of $3/7$ in place of $1/2$, but coming with a lower bound for the number of $n$’s satisfying $\|n\alpha\|<n^{-3/7+\varepsilon}$.
@@
They started in a similar way as we do here, converting the Diophantine condition $\|n\alpha\|<\delta$ into a condition of $n$ lying in a union of residue classes modulo $q$. Then they proceeded by detecting the squarefreeness of $n$ using the convolution identity $\mu^{2}(n)=\sum\limits_{t^{2}|n}\mu(t)$ and the congruence condition using the orthogonality relations for Dirichlet characters. After isolating a main term, they were left with character sums which were treated using Hölder’s inequality and moment bounds for character sums. This yielded an exponent of $1/2-\varepsilon$ for the relevant Diophantine approximation problem. To handle Diophantine approximation with denominators represented by sums of two squares following their approach, one may use the convolution identity $r_{2}(n)=4\sum\limits_{d|n}\chi_{4}(d)$ with $r_{2}(n)$ being the number of representations of $n$ as a sum of two squares and $\chi_{4}$ the non-trivial character modulo 4 . However, here we have a divisor condition of $d|n$ in place of $t^{2}|n$. As a consequence (even if one uses the hyperbola trick), one gets only an exponent of $1/3-\varepsilon$ instead of $1/2-\varepsilon$ in the Diophantine approximation problem for sums of two squares. We also note that their method does not extend to general quadratic forms $Q$ because of the lack of a simple formula for the number of representations of $n$ in the form $n=Q(x,y)$.
@@ 2. Preliminaries
where $A$ is an integral symmetric $2\times 2$ matrix with even diagonal elements, and $(x,y)^{T}$ is the column vector of the indeterminates. Then the adjoint matrix $A^{\dagger}$ of $A$ is defined by $A^{\dagger}A=\det(A)I_{2}$, where $\det(A)=\Delta$ is the determinant of the matrix $A$ and $I_{2}$ is the $2\times 2$ identity matrix. Clearly, $A^{\dagger}$ has also integer entries.
Our main tools are the Voronoi summation formula for quadratic forms and bounds for bilinear sums with Kloosterman sums to a fixed modulus. To state the Voronoi summation formula in this context, let us first set up some notations. For a positive integer $q$ and an integer $h$, we define the Gauss sum associated to the form $Q$ as
@@
The following lemma gives an evaluation of $G_{Q}(q,h)$ if $(2\Delta h,q)=1$.
@@
Suppose that $Q=\frac{1}{2}(x,y)A(x,y)^{T}$ is a PBQF with discriminant $\Delta$, and let $q,h$ be integers such that $(2\Delta h,q)=1$. Then we have
@@
$N(q,Q):=\sharp\{\vec{x}(\bmod q):A\vec{x}\equiv\vec{0}(\bmod q)\}$ (2.3)
and $\vec{x}=(x,y)^{T}.$
@@
See [3, Theorem 1] or [9, Theorem 1] for the particular case when $(2\Delta h,q)=1$. ∎
@@
For any $q\in\mathbb{N}$ and $h\in\mathbb{Z}$ with $(q,h)=1$, we have
@@
$B:=(b_{i}\delta_{ij})\quad\mbox{and}\quad A^{*}:=BV^{T}A^{\dagger}VB.$
@@
$\tilde{r}_{Q^{*}}(n,q,h):=q^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\ \sum_{\vec{y}(\bmod q)}e_{q}(hQ(\vec{y})).$ (2.4)
Now we are ready to state the Voronoi summation formula for the quadratic form $Q$.
@@
Let $f(n)$ be a continuously differentiable function on $\mathbb{R}$ with compact support in $\mathbb{R}_{\geq 0}$. Then under the conditions in Lemma 1, we have
$\sum_{n\in\mathbb{Z}}r_{Q}(n)e_{q}(nh)f(n)=2\pi\Delta^{-1/2}q^{-2}G_{Q}(q,h)\int\limits_{0}^{\infty}f(x)dx$
$+2\pi(q\sqrt{\Delta})^{-1}\sum_{n\in\mathbb{Z}}\tilde{r}_{Q^{*}}(n,q,h)e_{q}(-\overline{h}\overline{\Delta}n)\int\limits_{0}^{\infty}J_{0}\left(4\pi\frac{\sqrt{nx/\Delta}}{q}\right)f(x)dx,$
where $J_{0}(x)$ is the standard Bessel function and $\overline{\Delta}$ is a multiplicative inverse of $\Delta$ modulo $q$, i.e. $\overline{\Delta}\Delta\equiv 1\bmod{q}$.
@@
See the proof of [9, Equation (28)] for the special case when $(2\Delta h,q)=1$. ∎
@@
$S(m,n;q):=\sum_{\begin{subarray}{c}1\leq x\leq q-1\\ (x,q)=1\end{subarray}}e\left(\frac{mx+n\overline{x}}{q}\right).$
@@
The following Dirichlet approximation type lemma will allow us to choose the modulus $q$ coprime to $2\Delta$ (in fact coprime to any given non-zero integer).
@@
Fix $d\in\mathbb{N}$ and $\alpha\in\mathbb{R}\setminus\mathbb{Q}$. We aim to find good approximations of $\alpha$ by fractions $b/r$ with $(r,bd)=1$. Take one of the infinitely many approximations of $\alpha$ by $a/q$ with $(a,q)=1$ and